# Stock Price Responses to Firm-Level News in Supply Chain Networks

### v4

@@ Abstract
This study examines how positive and negative news about firms are associated with stock prices and whether these associations extend to suppliers and clients linked via supply chain relationships, using large samples of publicly listed firms worldwide and in Japan. News sentiment is measured using FinBERT, a natural language processing model fine-tuned for financial text, and supply chain links are identified from financial statements for global firms and from large-scale firm-level surveys for Japanese firms. We find that stock prices exhibit systematic associations with positive and negative news even before public disclosure. These associations are also observed for suppliers and clients before and after disclosure. In general, post-disclosure associations are larger than pre-disclosure associations, with the difference concentrated around the time of public news disclosure relative to the pre-disclosure period. However, for Japanese firms, the post-disclosure associations for suppliers and clients are smaller than the pre-disclosure associations, in contrast to the pattern observed for firms outside Japan.
@@ 1 Introduction
Firms interact with and influence each other through various types of networks. Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 23, 27] and knowledge networks through research collaboration [30, 29, 10, 72, 34]. Another type of network recently paid attention to in the academic literature and business and policy fields is supply chains formed through transactions of materials, parts, and components [15, 32]. One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other [6].
@@
Global supply chains have rapidly expanded because firms interconnected through supply chains can greatly benefit from the network. Most notably, supply chains can empower firms by enhancing the efficiency of production processes and thus optimizing firms’ economic performance [11, 5]. For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost-effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity [18]. Moreover, it is often evident that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services and from downstream to upstream firms because of learning of knowledge and technology [55, 72, 38].
@@
Conversely, supply chains can be a channel of negative shocks that aggravate firm performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability [9, 35, 7]. Such negative shocks as examined in the literature include the Great East Japan Earthquake in 2011 [49], the massive flood in Thailand in 2011 [31], the COVID-19 pandemic from 2020 to 2022 [17], and the Russia-Ukraine war that started in 2022 [54].
@@
Accordingly, prior studies show that positive or negative shocks propagate through supply chains, affecting the performance of firms linked through supply chains [7]. Because the corporate value of publicly listed firms is evaluated in the stock market [52, 71, 12], a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and client firms. For example, following the Boeing 737 MAX crashes in 2019 and 2020, stock prices of Boeing, some of its suppliers such as GE and Allegheny Technologies, and some of its clients such as American Airlines and Southwest Airlines declined sharply [13, 14, 22, 48, 62]. Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 was accompanied by stock price declines for Daihatsu, Toyota, and related firms including suppliers and clients such as Aisin [64, 63].
@@
These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains. Although relatively a lot of previous studies have examined the direct impact of shocks to firms on their own stock prices [44, 69, 68], there are not so many studies that have investigated the diffusion of shocks to other firms’ stock prices through supply chains. Cohen and Frazzini [16] show that information about a firm’s customers predicts future stock returns of its suppliers, providing early evidence that stock price reactions propagate through supply chain relationships. More recently, Le Tran and Coqueret [42] show that firm-level news shocks, particularly ESG-related news, affect not only the focal firm’s stock returns but also the returns of its suppliers and clients, providing direct evidence of news-driven spillovers along supply chain relationships. Complementing this evidence, Agarwal et al. [2] demonstrate that investor attention, measured through online co-search behavior, facilitates information diffusion across supply chain partners and contributes to predictable stock return comovements. Kim and Wagner [40] analyzed 315 cases reported by the Wall Street Journal regarding supply chain risks, which also included mentions of 69 suppliers and 246 clients. Pandit et al. utilized a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyze their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT [57]. In addition, similar studies focusing on China are discussed [45, 46]. However, to the best of our knowledge, no study has examined large-scale data spanning more than a decade and covering tens of thousands of firms.
To fill this gap, this study examines whether news sentiment is associated with stock price changes of firms and of their suppliers and clients in the pre- and post-disclosure periods, using two large firm-level samples, one for firms across the world and the other for Japanese firms. Specifically, we test whether the level of positive and negative sentiment about a firm in a news article affects the change rate of the stock price of the firm and its suppliers and clients before and after the disclosure of the news.
Beyond supply-chain-specific mechanisms, a growing literature examines how the content of news itself is quantified and linked to stock market reactions. Early studies quantified news content using dictionary-based and word-frequency approaches to capture investor sentiment and market reactions. For example, Tetlock [70] construct a sentiment measure based on word categories in financial news and show that media pessimism predicts short-run market returns and trading volume. Relatedly, Mizuno et al. [53] quantify the novelty and topicality of business news using word-level similarity measures and demonstrate that such text-based characteristics are strongly associated with subsequent stock market activity. In this field, neural networks are predominantly used. For example, hedge funds, such as Bridgewater Associates, utilize neural networks for predictions [8]. In academia, the transformer [73] is widely used [67, 25]. In addition to the transformer, there is a long history of the usage of neural networks to predict stock prices, such as long short-term memory [37, 41, 20, 21, 39, 43, 65, 26], convolutional neural networks [33], or other deep learning methods [47, 58, 66, 56]. We use sentiment analysis to determine the positiveness and negativeness of each news article [51, 28] and for this purpose, specifically, Bidirectional Encoder Representations from Transformers (BERT) [19] have garnered the most acclaim [69]. We take advantage of FinBERT [4], a fine-tuned model of BERT particularly for financial information.
@@
To advance this line of research from market- or firm-level settings to supply-chain networks and cross-country comparisons, it is essential to combine large-scale news data with comprehensive firm-level linkages and stock price information. We contribute to the literature by utilizing large-scale data in many respects. First, our firm-level data covers most publicly listed firms in the world and in Japan and identifies detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys. Second, our data on news articles are taken from the News Archive of Thomson Reuters and cover more than 20 million articles. Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive and negative news.
@@ 2 Data
2.1 News article
Our information on news relies on news articles written in English included in the News Archive provided by Thomson Reuters for the period 2003–2016. In the News Archive, multiple news articles may be filed for a single event. Typically, the first article is an alert, which is a short message containing only essential facts. This is followed by a newsbreak, usually released within 5–20 minutes after the alert. A newsbreak consists of a headline, which is similar to that of the alert, and two to four paragraphs of body text. Subsequent updates may follow the newsbreak after an additional 20–30 minutes, providing further information.
In this study, we use the newsbreak associated with each event. Although we are interested in the timing of public disclosure, alerts often do not include sufficient details about the related firms, whereas newsbreaks do. The time difference between alerts and newsbreaks is typically less than 20 minutes, which is negligible given that our analysis is conducted at the daily frequency. Accordingly, the release time of the newsbreak is aligned with the closest market closing time of the focal stock market, and the corresponding trading day is defined as day zero in the subsequent analyses. The newsbreak includes the time, headline, main text, Refinitiv Identification Codes (RICs) of one or more listed firms mentioned in the article, the language, and other metadata.
As we will explain in detail later, our sentiment analysis, which determines the degree of positivity or negativity of each news article about the firms mentioned, utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and its maximum is 3,911. Although the total number of news articles in the News Archive for the period examined is 20,803,561, we focus on 3,447,425 that mention any firm publicly listed in 105 stock markets in the world. Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33 (Figure 1). The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed. 55.6% of firms were never mentioned, while 0.1% were mentioned more than 10,000 times (Figure 2).
@@ 2.1 News article
Stock price
We utilize daily stock prices of currently listed firms in global stock markets, obtained from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, through its Application Programming Interface (API). Firms that were listed in the past but are no longer listed are excluded from the sample, as their stock price data are not available through the API. We define the daily stock price of each firm as its closing price. Accordingly, the stock price data include Refinitiv Identification Codes (RICs), dates, and closing prices.
In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend. The Refinitiv index is available for 105 major markets that cover 92% of listed firms worldwide.
Supply chain
To identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003–2016 obtained as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statements. The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [24]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and websites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed recently and thus ignores firms in the distant past. In a later section, the recorded supply chain links at the time are used to examine news-related stock price changes along supply chain relationships.
The FactSet supply chain database does not cover the universe of firms worldwide. Instead, it identifies major customer–supplier relationships disclosed in financial statements and related sources, which are typically limited to economically significant partners. As a result, its coverage is intentionally limited, especially in earlier years.
FactSet uses its own identifier system, referred to as the FactSet Security Identifier. We convert these identifiers to Refinitiv Identification Codes (RICs) in order to link the FactSet data with news and stock price data.
@@
To identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms for each surveyed firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Since those firms are relatively large, their partners report them as important suppliers or clients. Therefore, although the numbers of suppliers and clients are limited to 23 in the reports, large firms can have many partners. Accordingly, the maximum number of suppliers and clients is 12,729 and 12,016, respectively, in 2016, providing broad coverage of supply chains of Japanese firms.
The TSR database provides exceptionally broad coverage of firm-level relationships, including both listed and unlisted firms. In this study, however, we restrict our analysis to relationships involving listed firms and links between such firms, in order to ensure consistency with market-based information. Listed firms in the TSR database are identified by corporate IDs and converted into Refinitiv Identification Codes (RICs). The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
@@ 3 Methods
Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT [4]. FinBERT is a variant of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018) [19]. Specifically, FinBERT uses Reuters’ TRC2-financial, which consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank, which consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three sentiments about the firms mentioned in the article. Examples are shown in Table 3. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end. The distribution of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.

### v5

@@ Abstract
This study examines how positive and negative news about firms is associated with stock prices and whether these associations extend to firms’ suppliers and clients connected through supply chain relationships, using large samples of publicly listed firms worldwide and in Japan. News sentiment is measured using FinBERT, a natural language processing model fine-tuned for financial texts, and supply chain links are identified from financial statements for global firms and from large-scale firm-level surveys for Japanese firms. We find that stock prices exhibit systematic associations with positive and negative news even before public disclosure. These associations are also observed for suppliers and clients before and after disclosure. In general, relative to the pre-disclosure period, post-disclosure associations are larger, with the difference concentrated around the time of public news disclosure. However, for Japanese firms, the post-disclosure associations for suppliers and clients are smaller than the pre-disclosure associations, in contrast to the pattern observed for firms outside Japan.
@@ 1 Introduction
Firms interact and influence one another through various types of networks. Major firm networks that play a significant role in firms’ behavior and performance in the literature include financial networks through ownership [2, 24, 27] and knowledge networks through research collaboration [30, 29, 11, 72, 34]. Another type of network that has received increasing attention in the academic literature and business and policy fields is supply chains formed through transactions in materials, parts, and components [16, 32]. One reason for the growing attention is that supply chains have expanded globally, linking firms worldwide within a small number of steps to one another [7].
@@
Global supply chains have rapidly expanded because firms interconnected through supply chains can greatly benefit from the network. Most notably, supply chains can empower firms by enhancing the efficiency of production processes and thus improving firms’ economic performance [12, 6]. For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost-effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity [19]. Moreover, it is frequently observed that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services, and from downstream to upstream firms because of the learning and diffusion of knowledge and technology [55, 72, 38].
@@
Conversely, supply chains can be a channel of negative shocks that adversely affect firms’ performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability [10, 35, 8]. Such negative shocks examined in the literature include the Great East Japan Earthquake in 2011 [49], the massive flood in Thailand in 2011 [31], the COVID-19 pandemic from 2020 to 2022 [18], and the Russia-Ukraine war that started in 2022 [54].
@@
Accordingly, prior studies show that positive or negative shocks propagate through supply chains, affecting the performance of firms linked through supply chains [8]. Because the corporate value of publicly listed firms is assessed in the stock market [52, 71, 13], a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and clients. For example, following the Boeing 737 MAX crashes in 2019 and 2020, the stock prices of Boeing, some of its suppliers such as GE and Allegheny Technologies, and some of its clients such as American Airlines and Southwest Airlines declined sharply [15, 14, 23, 48, 62]. Also, misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 was accompanied by stock price declines for Daihatsu, Toyota, and related firms including suppliers and clients such as Aisin [64, 63].
@@
These examples show that shocks can propagate through supply chains and affect the stock prices of firms connected through supply chains. Although a substantial body of previous studies has examined the direct impact of shocks to firms on their own stock prices [44, 69, 68], there are relatively few studies that have investigated the diffusion of shocks to other firms’ stock prices through supply chains. Cohen and Frazzini [17] show that information about a firm’s customers predicts future stock returns of its suppliers, providing early evidence that stock price reactions propagate through supply chain relationships. More recently, Le Tran and Coqueret [42] show that firm-level news shocks, particularly ESG-related news, affect not only the focal firm’s stock returns but also the returns of its suppliers and clients, providing direct evidence of news-driven spillovers along supply chain relationships. Complementing this evidence, Agarwal et al. [3] demonstrate that investor attention, measured through online co-search behavior, facilitates information diffusion across supply chain partners and contributes to predictable stock return comovements. Kim and Wagner [40] analyzed 315 cases reported by the Wall Street Journal regarding supply chain risks, which also included mentions of 69 suppliers and 246 clients. Pandit et al. used a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyzed their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT [57]. In addition, similar studies focusing on China are discussed [45, 46]. However, existing studies remain constrained in two important respects. First, most analyses focus on stock price reactions observed after news disclosures or identifiable events, leaving stock price movements prior to disclosure largely unexplored. Second, evidence based on large-scale data covering long time horizons and extensive supply-chain networks remains scarce.
@@
To fill this gap, this study examines whether news sentiment is associated with stock price changes of firms and their suppliers and clients in the pre- and post-disclosure periods, extending the analysis from market- or firm-level settings to supply-chain networks and cross-country comparisons. Using two large firm-level samples, one covering most publicly listed firms worldwide and the other focusing on Japanese firms, we test whether the level of positive and negative sentiment about a firm in a news article is associated with changes in the stock prices of the firm and its suppliers and clients before and after news disclosure. Our analysis combines more than 20 million news articles from the Thomson Reuters News Archive, comprehensive firm-level linkages that identify detailed supply chains based on financial statements, news articles, websites, and firm-level surveys, and daily stock price data for all listed firms, enabling an examination of both short- and long-run associations.
In this study, we adopt a reduced-form perspective and examine the relationship between news sentiment and stock prices without explicitly specifying investors’ information processing or decision-making. This abstraction includes how investors perceive and incorporate supply-chain relationships into their trading behavior.
@@ 2 Data
2.1 News Articles
We use news articles written in English from the Thomson Reuters News Archive for the period 2003–2016. In the News Archive, multiple news articles may be filed for a single event. Typically, the first article is an alert, which is a short message containing only essential facts. This is followed by a newsbreak, usually released within 5–20 minutes after the alert. A newsbreak consists of a headline, which is similar to that of the alert, and two to four paragraphs of body text. Subsequent updates may follow the newsbreak after an additional 20–30 minutes, providing further information.
In this study, we use the newsbreak corresponding to each event. Although we are interested in the timing of public disclosure, alerts often do not include sufficient details about the related firms, whereas newsbreaks do. The time difference between alerts and newsbreaks is typically less than 20 minutes, which is negligible given that our analysis is conducted at the daily frequency. Accordingly, the release time of the newsbreak is aligned with the next market closing time of the relevant stock market, and the corresponding trading day is defined as day zero in the subsequent analyses. The newsbreak includes the time, headline, main text, Refinitiv Identification Codes (RICs) of one or more listed firms mentioned in the article, and other metadata.
As we will explain in detail later, our sentiment analysis, which determines the degree of positivity or negativity of each news article, utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and its maximum is 3,911. Although the total number of news articles in the News Archive for the period examined is 20,803,561, we focus on 3,447,425 that mention any publicly listed firm in 105 stock markets in the world. Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33 (Figure 1). The number of news articles that mention a particular listed firm during the period 2003–2016 is highly skewed. More than half of firms (55.6%) were never mentioned, while 0.1% were mentioned more than 10,000 times (Figure 2).
@@ 2.1 News Articles
Stock Prices
We utilize daily stock prices of firms that are currently listed in global stock markets, obtained from the Refinitiv Eikon database, one of the largest providers of financial markets data and infrastructure, through its Application Programming Interface (API). Firms that were listed in the past but are no longer listed are excluded from the sample, as their stock price data are not available through the API. We define the daily stock price of each firm as its closing price. Accordingly, the stock price data include RICs, dates, and closing prices.
In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average stock price level in the market to control for the market-specific time trend. The Refinitiv index is available for 105 major markets that cover 92% of listed firms worldwide.
Supply Chains
To identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms worldwide for the period 2003–2016, as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statements. The US, Japanese, and many other governments require every listed firm to disclose its major clients whose sales account for more than 10% of its total sales, at least in the period examined in this study [1]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and websites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed recently and thus does not cover firms in the distant past. In a later section, the recorded supply chain links at the time are used to examine news-related stock price changes along supply chain relationships.
The FactSet supply chain database does not cover the full universe of firms worldwide. Instead, it identifies major customer–supplier relationships disclosed in financial statements and related sources, which are typically limited to economically significant partners. As a result, its coverage is intentionally limited, especially in earlier years.
FactSet uses its own identifier system, referred to as the FactSet Security Identifier. We convert these identifiers to RICs in order to link the FactSet data with news and stock price data.
@@
To identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms for each surveyed firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Since those firms are relatively large, their partners report them as important suppliers or clients. Therefore, although the numbers of suppliers and clients are limited to 23 in the reports, large firms can have many partners. Accordingly, the maximum number of suppliers and clients is 12,729 and 12,016, respectively, in 2016, thus providing broad coverage of supply chains of Japanese firms.
The TSR database provides exceptionally broad coverage of firm-level relationships, including both listed and unlisted firms. In this study, however, we restrict our analysis to relationships involving listed firms and links between such firms, in order to ensure consistency with market-based information. Listed firms in the TSR database are identified by corporate IDs and converted into RICs. The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
@@ 3 Methods
A growing literature examines how news content itself is quantified and linked to stock market reactions. Early studies quantified news content using dictionary-based and word-frequency approaches to capture investor sentiment and market reactions. For example, Tetlock [70] construct a sentiment measure based on word categories in financial news and show that media pessimism predicts short-run market returns and trading volume. Relatedly, Mizuno et al. [53] quantify the novelty and topicality of business news using word-level similarity measures and demonstrate that such text-based characteristics are strongly associated with subsequent stock market activity. In this field, neural networks are predominantly used. For example, hedge funds, such as Bridgewater Associates, utilize neural networks for predictions [9]. In academia, the transformer [73] is widely used [67, 25]. In addition to the transformer, there is a long history of using neural networks to predict stock prices, such as long short-term memory [37, 41, 21, 22, 39, 43, 65, 26], convolutional neural networks [33], or other deep learning methods [47, 58, 66, 56]. We use sentiment analysis to determine the positiveness and negativeness of each news article [51, 28] and for this purpose, in particular, Bidirectional Encoder Representations from Transformers (BERT) [20] have garnered the most acclaim [69]. We take advantage of FinBERT [5], a fine-tuned model of BERT particularly for financial information.
Our sentiment analysis to determine the degree of positive and negative sentiment of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT [5]. FinBERT is a variant of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018) [20]. Specifically, FinBERT uses Reuters’ TRC2-financial, which consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank, which consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three sentiments about the firms mentioned in the article. Examples are shown in Table 3. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, International Securities Identification Numbers, and fixed phrases at the end. The distributions of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.